Solve the initial value problem below using the method of Laplace transforms. y +4y= 1662 - 12t + 16, y(0) = 0, y(O) = 7 C
Answer #1

y + 4y 16-12t +16 4(o)=0 y(o) 27 Taking Laplace Teans form on both sides <[y +4Y] = L[16+²-12t+16 2[yil] +42[y] = 162{+2]Taking Partial fractions 753 +1652 - 125+32 OSTE + S sa s3 (374) 2 +(ostels 75 +1657-125 +32 zestur A (5²) (5²74) + B (3) Cs- 22 - 2/3 -25+10 shtu. Ycs) s 3 Applying - 2s +10 2- 3 + st s3 stu YCE) : inveese Caplace Transform on both sides <-{xes)] =